# RoPINN: Region Optimized Physics-Informed Neural Networks

### v1

@@ Abstract
Physics-informed neural networks (PINNs) have been widely applied to solve partial differential equations (PDEs) by enforcing outputs and gradients of deep models to satisfy target equations. Due to the limitation of numerical computation, PINNs are conventionally optimized on finite selected points. However, since PDEs are usually defined on continuous domains, solely optimizing models on scattered points may be insufficient to obtain an accurate solution for the whole domain. To mitigate this inherent deficiency of the default scatter-point optimization, this paper proposes and theoretically studies a new training paradigm as region optimization. Concretely, we propose to extend the optimization process of PINNs from isolated points to their continuous neighborhood regions, which can theoretically decrease the generalization error, especially for hidden high-order constraints of PDEs. A practical training algorithm, Region Optimized PINN (RoPINN), is seamlessly derived from this new paradigm, which is implemented by a straightforward but effective Monte Carlo sampling method. By calibrating the sampling process into trust regions, RoPINN finely balances sampling efficiency and generalization error. Experimentally, RoPINN consistently boosts the performance of diverse PINNs on a wide range of PDEs without extra backpropagation or gradient calculation.
@@ 3.1 Region Optimization
We discuss the generalization error in expectation [13], which is independent of the point selection, thereby quantifying the error of PINN optimization more rigorously.
@@
Suppose that the loss function $\mathcal{L}$ is $L$-Lipschitz-$\beta$-smooth for $\theta$, if we run stochastic gradient method with step size $\alpha_{t}$ at the $t$-th step for $T$ iterations, we have that:
@@
(2) If $\mathcal{L}$ is bounded by a constant $C$ for all $\theta,\boldsymbol{x}$ and is non-convex for $\theta$ with monotonically non-increasing step sizes $\alpha_{t}\leq\frac{1}{\beta t}$, then $\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}$ (tighter bound than [47]).
@@
Suppose that the point optimization loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth for $\theta$, if we run stochastic gradient method with step size $\alpha_{t}$ for $T$ iterations based on region optimization loss $\mathcal{L}_{r}^{\mathrm{region}}$ in Eq. (5), the generalization error in expectation satisfies:
(1) If $\mathcal{L}$ is convex for $\theta$ and $\alpha_{t}\leq\frac{2}{\beta}$, then $\mathcal{E}_{\mathrm{gen}}\leq(1-\frac{|\Omega_{r}|^{2}}{2^{(d+1)}|\Omega|})\frac{2L^{2}}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t}$.
(2) If $\mathcal{L}$ is bounded by a constant $C$ and is non-convex for $\theta$ with monotonically non-increasing step sizes $\alpha_{t}\leq\frac{1}{\beta t}$, then $\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-JL(\frac{|\Omega_{r}|^{2}}{2^{(d+1)}|\Omega|})^{2}$, where $J$ is a finite number that depends on the training property at the several beginning iterations.
From Theorems 3.3 and 3.5, we can observe that region optimization can reduce the generalization error $\mathcal{E}_{\mathrm{gen}}$. Note that in most cases of PDEs, $d\leq 3$. Larger $\Omega_{r}$ will facilitate better generalizability. Our method will degenerate to point optimization when $\Omega_{r}=0$, in which a single point is selected.
@@
In our proposed region optimization (Eq. (5)), the integral operation on the input domain can also relax the smoothness requirements of the loss function $\mathcal{L}$. For example, without any assumption of the smoothness of $\mathcal{L}(u_{\theta},\boldsymbol{x})$ on $\boldsymbol{x}$, we can directly derive the generalization error for the first-order loss $\frac{\partial}{\partial x_{j}}\mathcal{L}_{r}^{\mathrm{region}}(u_{\theta},\boldsymbol{x})$ on the $j$-th dimension as follows.
@@
Suppose that $\mathcal{L}$ is bounded by $C$ for all $\theta,\boldsymbol{x}$ and is $L$-Lipschitz and $\beta$-smooth for $\theta$, if we run stochastic gradient method based on first-order $j$-th dimension loss function $\frac{\partial}{\partial x_{j}}\mathcal{L}_{r}^{\mathrm{region}}$ for $T$ iterations, the generalization error in Theorem 3.5(2) still holds when we adopt the monotonically non-increasing step size $\alpha_{t}\leq\frac{1}{2\beta t}$.
Corollary 3.6 implies that the integral on the input domain in region optimization can help training PINNs with high-order constraints, which is valuable for high-order PDEs, such as wave equations.
@@ 3.2 Practical Algorithm
$\begin{split}\mathbb{E}_{\boldsymbol{\xi}\sim{U}[0,r]^{d+1}}\big{(}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})\big{)}=\mathbb{E}_{\boldsymbol{\xi}\sim{U}[0,r]^{d+1}}\left(\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x})+\boldsymbol{\xi}\nabla_{\theta}\mathcal{L}_{1}(u_{\theta},\boldsymbol{x})+\mathcal{O}(\boldsymbol{\xi}^{2})\right),\end{split}$ (9)
@@
Suppose that there exists a constant $H$, s.t. $\forall\boldsymbol{v}$ and $\forall\boldsymbol{x}\in\Omega$, $\left|\boldsymbol{v}^{\sf T}\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}(u_{\theta},\boldsymbol{x})\boldsymbol{v}\right|\leq H\|\boldsymbol{v}\|^{2}$, if the step size $\alpha_{t}=\frac{1}{\sqrt{t+1}}$ decreases over time for $T$ iterations, the region optimization based on Monte Carlo approximation will converge at the speed of
@@
Theorem 3.9 (Estimation error).
@@
Suppose that loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth for $\theta$ and the $t$-th step parameter is $\theta_{t}$, the gradient difference between successive iterations $\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\|$ will converge to $\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})\|$, which is bounded by the following inequality:
$\begin{split}\left\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\right\|-\left\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})\right\|\leq\beta L\alpha_{t-1},\end{split}$ (13)
where $\alpha_{t-1}$ represents the step size at the $(t-1)$-th iteration.
@@
$\begin{split}\lim_{t\to\infty}\sigma\left({\nabla_{\theta}\mathcal{L}(u_{\theta_{t-i+1}},\boldsymbol{z}_{i})}_{i=1}^{{T_{0}}}\right)=\sigma\left({\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{i})}_{i=1}^{{T_{0}}}\right).\end{split}$ (14)
@@
Recall that in Theorem 3.5, we observe that a larger region size will benefit the generalization error, while Theorem 3.9 demonstrates that too large region size will also cause unstable training because it will result in excessive estimation error of Monte Carlo sampling in our implementation. Thus, our design in calibrating trust region can adaptively balance the generalization error and training stability.
@@ 4 Experiments
In RoPINN (Algorithm 1), we set the multi-iteration hyperparameter ${T_{0}}=5$ and the initial region size $r=10^{-4}$ for all datasets, where the trust region size will be adaptively adjusted to fit the PDE property during training. For 1D-Reaction, 1D-Wave and Convection, we follow [53] and train the model with L-BFGS optimizer [25] for 1,000 iterations. As for PINNacle, we strictly follow their official configuration [12] and train the model with Adam [20] for 20,000 iterations. Besides, for simplicity and fair comparison, we set the weights of PINN loss as equal, that is $\lambda_{\ast}=1$ in Eq. (2). Canonical loss formalized in Eq. (2), relative L1 error (rMAE) and relative L2 error (rMSE) are recorded. All experiments are implemented in PyTorch [31] and trained on a single NVIDIA A100 GPU. See Appendix C for more implementation details.
@@ 4.2 Algorithm Analysis
For efficiency, RoPINN only samples one point within the trust region to approximate the region gradient descent. However, it is worth noticing that sampling more points will make the approximation in Eq. (8) more accurate, which can further make the optimization process adapt to a larger region size $r$. Thus, as shown in Figure 2(a-b), sampling more points will also increase the finally learned region size $r$ and speed up the convergence. In addition, Figure 2(c) shows that adding sampled points can also improve the final performance, which has been proved in Theorem 3.5 that the upper bound of generalization error is proportion to region size.
@@
To verify the effectiveness of our design in RoPINN, we present ablations in Figure 4. It is observed that although we only sample one point, even fixed-size region optimization can also boost the performance of PINNs in most cases. However, as illustrated in Theorem 3.9, the sampling process may also cause estimation error, so the relative promotion is inconsistent and unstable among different PDEs and base models. With our proposed trust region calibration, we can obtain a more significant and consistent improvement, verifying the effectiveness of our design in RoPINN.
@@
Figure 5: Loss landscape of RoPINN and vanilla PINNs on Convection. Error refers to $(u_{\theta}-u)$.
@@ 5 Conclusion
This paper presents and analyzes a new PINN optimization paradigm: region optimization. Going beyond previous scatter-point optimization, we extend the optimization from selected points to their neighborhood regions. Based on this idea, RoPINN is implemented as a simple but effective training algorithm, where an effcient Monte Carlo approximation process is used along with a trust region calibration strategy to control the estimation error. In addition to theoretical advantages, RoPINN can consistently boost the performance of various advanced PINN models without extra backpropagation or gradient calculation, demonstrating favorable efficiency, training stability and general capability.
@@
$\begin{split}\mathbb{E}_{\mathcal{S},\mathcal{A}}\left[\mathcal{L}(\mathcal{A}(\mathcal{S}),\mathcal{S})\right]&=\mathbb{E}_{\mathcal{S},\mathcal{A}}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\mathcal{A}(\mathcal{S}),\boldsymbol{x}_{i})\right]\\ &=\mathbb{E}_{\mathcal{S},\mathcal{S}^{\prime},\mathcal{A}}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\mathcal{A}(\mathcal{S}^{(i)}),\boldsymbol{x}_{i}^{\prime})\right]\quad\quad\quad\quad\text{(Expectation on $\boldsymbol{x}_{i}^{\prime}$)}\\ &=\mathbb{E}_{\mathcal{S},\mathcal{S}^{\prime},\mathcal{A}}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\mathcal{A}(\mathcal{S}),\boldsymbol{x}_{i}^{\prime})\right]+\delta\\ &=\mathbb{E}_{\mathcal{S},\mathcal{A}}\left[\mathcal{L}(\mathcal{A}(\mathcal{S}),\Omega)\right].\end{split}$ (16)
@@
According to Lemma A.1, we attempt to bound the generalization error in expectation $\mathcal{E}_{\mathrm{gen}}$ by analyzing the error difference between two selected sample sets. Denote that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are identical except for one point. Suppose that with the stochastic gradient method on these two sets, we can obtain two optimization trajectories $\{\theta_{t}\}_{t=1}^{T}$ and $\{\theta_{t}^{\prime}\}_{t=1}^{T}$ respectively. According to Assumption 3.2, we have the following inequality:
@@
Note that at the $t$-th step, with probability $(1-\frac{1}{|\mathcal{S}|})$, the example selected by the stochastic gradient method is the same in both $\mathcal{S}$ and $\mathcal{S}^{\prime}$. As for the other $\frac{1}{|\mathcal{S}|}$ probability, we have to deal with different selected samples. Thus, according to Lemma A.2, we have:
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|\right]&=(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[G_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G_{\alpha_{t},\boldsymbol{x}}(\theta_{t}^{\prime})\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[G_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})\mathbb{E}[\|\theta_{t}-\theta_{t}^{\prime}\|]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+\|\alpha_{t}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x})-\alpha_{t}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}^{\prime})\|\right].\\ \end{split}$ (20)
@@
We also consider the optimization trajectory $\{\theta_{t}\}_{t=1}^{T}$ and $\{\theta_{t}^{\prime}\}_{t=1}^{T}$ from $\mathcal{S}$ and $\mathcal{S}^{\prime}$, which are identical except for one element. Let $\delta_{t}=\|\theta_{t}-\theta_{t}^{\prime}\|$ and $t_{0}\in\{1,\cdots,|\mathcal{S}|\}$ be a considered iteration, then we have:
@@
Firstly, we would like to prove the convex case as follows.

### v2

@@ Abstract
Physics-informed neural networks (PINNs) have been widely applied to solve partial differential equations (PDEs) by enforcing outputs and gradients of deep models to satisfy target equations. Due to the limitation of numerical computation, PINNs are conventionally optimized on finite selected points. However, since PDEs are usually defined on continuous domains, solely optimizing models on scattered points may be insufficient to obtain an accurate solution for the whole domain. To mitigate this inherent deficiency of the default scatter-point optimization, this paper proposes and theoretically studies a new training paradigm as region optimization. Concretely, we propose to extend the optimization process of PINNs from isolated points to their continuous neighborhood regions, which can theoretically decrease the generalization error, especially for hidden high-order constraints of PDEs. A practical training algorithm, Region Optimized PINN (RoPINN), is seamlessly derived from this new paradigm, which is implemented by a straightforward but effective Monte Carlo sampling method. By calibrating the sampling process into trust regions, RoPINN finely balances optimization and generalization error. Experimentally, RoPINN consistently boosts the performance of diverse PINNs on a wide range of PDEs without extra backpropagation or gradient calculation. Code is available at this repository: https://github.com/thuml/RoPINN.
@@ 3.1 Region Optimization
Here we discuss the generalization error in expectation [13], which is independent of the point selection, thereby quantifying the error of PINN optimization more rigorously.
@@
Suppose that the loss function $\mathcal{L}$ is $L$-Lipschitz-$\beta$-smooth for $\theta$. If we run stochastic gradient method with step size $\alpha_{t}$ at the $t$-th step for $T$ iterations, we have that:
@@
(2) If $\mathcal{L}$ is bounded by a constant $C$ for all $\theta,\boldsymbol{x}$ and is non-convex for $\theta$ with monotonically non-increasing step sizes $\alpha_{t}\leq\frac{1}{\beta t}$, then $\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}$ (tighter bound than [13, 52]).
@@
Suppose that the point optimization loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth for $\theta$. If we run stochastic gradient method with step size $\alpha_{t}$ for $T$ iterations based on region optimization loss $\mathcal{L}_{r}^{\mathrm{region}}$ in Eq. (5), the generalization error in expectation satisfies:
(1) If $\mathcal{L}$ is convex for $\theta$ and $\alpha_{t}\leq\frac{2}{\beta}$, then $\mathcal{E}_{\mathrm{gen}}\leq(1-\frac{|\Omega_{r}|}{|\Omega|})\frac{2L^{2}}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t}$.
(2) If $\mathcal{L}$ is bounded by a constant $C$ for all $\theta,\boldsymbol{x}$ and is non-convex for $\theta$ with monotonically non-increasing step sizes $\alpha_{t}\leq\frac{1}{\beta t}$, then $\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-JL(\frac{|\Omega_{r}|}{|\Omega|})^{2}$, where $J$ is a finite number that depends on the training property at the several beginning iterations.
Proof.
Based on the Lipschitz assumption, $\mathcal{E}_{\mathrm{gen}}$ can be bounded by times the expectation of distance between parameter $\theta$ optimized from different training sets. Region optimization paradigm will bring a more “consistent” gradient optimization direction than point optimization at each iteration, thereby benefitting the generalization property. See Appendix A.3 for complete proof. ∎
From Theorems 3.3 and 3.5, we can observe that region optimization can reduce the generalization error $\mathcal{E}_{\mathrm{gen}}$. Furthermore, the region optimization theorem also provides a more general theoretical framework. For example, the conventional point optimization is equivalent to the case of $\Omega_{r}=0$, where only one single point is selected for each region. For another extreme case, enlarging the region size to the whole domain (i.e. $\Omega_{r}=\Omega$), Eq. (5) is equivalent to directly optimizing the loss defined on $\Omega$, where the generalization error will be reduced to zero. Unfortunately, this ideal situation cannot be satisfied in practice, since Eq. (5) requires the precise calculation of the integral of the whole domain. More discussions about practical implementation are deferred to the next section.
@@
In our proposed region optimization (Eq. (5)), the integral operation on the input domain can also relax the smoothness requirements of the loss function $\mathcal{L}$. For example, without any additional assumption of the smoothness of $\mathcal{L}(u_{\theta},\boldsymbol{x})$ on $\boldsymbol{x}$, we can directly derive the generalization error for the first-order loss $\frac{\partial}{\partial x_{j}}\mathcal{L}_{r}^{\mathrm{region}}(u_{\theta},\boldsymbol{x})$ on the $j$-th dimension as follows.
@@
Suppose that $\mathcal{L}$ is bounded by $C$ for all $\theta,\boldsymbol{x}$ and is $L$-Lipschitz and $\beta$-smooth for $\theta$. If we run stochastic gradient method based on first-order $j$-th dimension loss function $\frac{\partial}{\partial x_{j}}\mathcal{L}_{r}^{\mathrm{region}}$ for $T$ iterations, the generalization error in Theorem 3.5(2) still holds when we adopt the monotonically non-increasing step size $\alpha_{t}\leq\frac{1}{2\beta t}$.
Corollary 3.6 implies that the integral on the input domain in region optimization can help training PINNs with high-order constraints, which is valuable for high-order PDEs, such as wave equations. In contrast, this valuable property cannot be achieved in the classic point optimization. See Example 3.7.
@@ 3.2 Practical Algorithm
$\begin{split}\mathbb{E}_{\boldsymbol{\xi}\sim{U(\Omega_{r})}}\big{(}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})\big{)}=\mathbb{E}_{\boldsymbol{\xi}{\sim U(\Omega_{r})}}\left(\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x})+{\nabla_{\theta}(\boldsymbol{\xi}^{\sf T}\mathcal{L}_{1}}(u_{\theta},\boldsymbol{x}))+\mathcal{O}({\|\boldsymbol{\xi}\|^{2}})\right),\end{split}$ (9)
@@
Suppose that there exists a constant $H$, s.t. $\forall\boldsymbol{v}$ and $\forall\boldsymbol{x}\in\Omega$, $\left|\boldsymbol{v}^{\sf T}\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}(u_{\theta},\boldsymbol{x})\boldsymbol{v}\right|\leq H\|\boldsymbol{v}\|^{2}$. If the step size $\alpha_{t}=\frac{1}{\sqrt{t+1}}$ decreases over time for $T$ iterations, the region optimization based on Monte Carlo approximation will converge at the speed of
@@
Theorem 3.9 (Gradient estimation error).
@@
Suppose that loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth for $\theta$ and the $t$-th step parameter is $\theta_{t}$. Two gradient difference sequences between successive iterations, $\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\|$ and $\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})\|$, share the same limit, as the difference of the two sequences is dominated by the following inequality:
$\begin{split}{\big{|}}\left\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\right\|-\left\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})\right\|{\big{|}}\leq\beta L\alpha_{t-1},\end{split}$ (13)
where $\alpha_{t-1}$ represents the step size at the $(t-1)$-th iteration, which approaches 0 as $t$ tends to $\infty$.
@@
$\begin{split}\lim_{t\to\infty}\sigma\left(\left\{\nabla_{\theta}\mathcal{L}(u_{\theta_{t-i+1}},\boldsymbol{z}_{i})\right\}_{i=1}^{{T_{0}}}\right)=\sigma\left(\left\{\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{i})\right\}_{i=1}^{{T_{0}}}\right).\end{split}$ (14)
@@
Balance of Generalization and Optimization
Recall that in Theorem 3.5, we observe that a larger region size will benefit the generalization error, while Theorem 3.9 demonstrates that too large region size will also cause unstable training because it will result in excessive gradient estimation error of Monte Carlo sampling in our implementation. The above analysis reveals the underlying trade-off between generalization and optimization of PINN models, which is concretely formalized as follows.
Theorem 3.12 (Region Optimization with gradient estimation error).
Based on the same assumption in Theorem 3.5 but optimize the model with the approximated region optimization loss $\mathcal{L}_{r}^{\mathrm{approx}}(u_{\theta},\boldsymbol{x})=\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi}),\boldsymbol{\xi}\sim U(\Omega_{r})$ for $T$ iterations, we further denote the upper bound of gradient estimation error as ${{\mathcal{E}_{r,\mathrm{grad}}}}=\max_{t\leq T}\|\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{approx}}-\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}\|$, then $\mathcal{E}_{\mathrm{gen}}$ satisfies:
[…]
Proof.
Compared to Theorem 3.5, considering the gradient estimation error will bring extra optimization discrepancy between different training sets. See Appendix B.4 for complete proof. ∎
Based on Theorem 3.12, we can find that classical point optimization ($\Omega_{r}=0$) and sampling points globally ($\Omega_{r}=\Omega$) discussed in Theorem 3.5 correspond to two extreme cases. The former makes ${\mathcal{E}_{r,\mathrm{grad}}}=0$ but cannot reduce the generalization error, while the latter one holds a large gradient estimation error. Thus, neither case can bring a perfect generalization. In contrast, the design for calibrating trust regions in RoPINN provides an adaptive strategy to better balance generalization and optimization, which can adjust the region size according to multi-iteration training stability.
@@ 4 Experiments
In RoPINN (Algorithm 1), we select the multi-iteration hyperparameter ${T_{0}}$ from $\{5,10\}$ and set the initial region size $r=10^{-4}$ for all datasets, where the trust region size will be adaptively adjusted to fit the PDE property during training. For 1D-Reaction, 1D-Wave and Convection, we follow [58] and train the model with L-BFGS optimizer [27] for 1,000 iterations. As for PINNacle, we strictly follow their official configuration [12] and train the model with Adam [21] for 20,000 iterations. Besides, for simplicity and fair comparison, we set the weights of PINN loss as equal, that is $\lambda_{\ast}=1$ in Eq. (2). Canonical loss formalized in Eq. (2), relative L1 error (rMAE) and relative L2 error (rMSE) are recorded. All experiments are implemented in PyTorch [34] and trained on a single NVIDIA A100 GPU. See Appendix C for more implementation details.
@@ 4.2 Algorithm Analysis
For efficiency, RoPINN only samples one point within the trust region to approximate the region gradient descent. However, it is worth noticing that sampling more points will make the approximation in Eq. (8) more accurate, leading to a lower gradient estimation error. Further, since RoPINN employs an adaptive strategy to adjust region $\Omega_{r}$, a lower gradient estimation error will also make the optimization process adapt to a larger region size $r$. Therefore, we observe in Figure 3(a-b) that sampling more points will also increase the finally learned region size $r$ and speed up the convergence. In addition, Figure 3(c) shows that adding sampled points can also improve the final performance, which has also been theoretically justified in Theorem 3.12 that the upper bound of generalization error is inversely proportion to gradient estimation error.
@@
Efficiency analysis
As we discussed above, sampling more points can benefit the final performance, while we choose only sample one point as the default setting of RoPINN in the spirit of boosting PINNs without extra backpropagation or gradient calculation, which has already achieved significant promotion w.r.t. original PINNs (Table 2). To provide a more comprehensive understanding of algorithm property, we plot the efficiency-performance curve in Figure 4, where we can obtain the following observations. Firstly, computation costs will grow linearly when adding points. Secondly, more points will bring better performance but will saturate around 10 points, where the performance fluctuations of 9, 13, and 30 points are within three times the standard deviations (Appendix D.3).
Figure 4: Efficiency and model performance w.r.t. number of samples. Note that the default setting of RoPINN is just sampling one point, which will not bring extra gradient calculation costs.
@@
To verify the effectiveness of our design in RoPINN, we present ablations in Figure 5. It is observed that although we only sample one point, even fixed-size region optimization can also boost the performance of PINNs in most cases, demonstrating the effectiveness of introducing “region” to PINN optimization. However, as illustrated in Theorem 3.9, the sampling process may also cause gradient estimation error, so the relative promotion is inconsistent and unstable among different PDEs and base models. With our proposed trust region calibration, we can obtain a more significant and consistent improvement, indicating that achieving a better balance between optimization and generalization (formalized in Theorem 3.12) performs an essential role in training PINN models.
@@
Figure 6: Loss landscape of RoPINN and vanilla PINNs on the Convection equation. Error Map refers to the distance between model prediction and the accurate solution, i.e. $(u_{\theta}-u)$.
@@ 5 Conclusion
This paper presents and analyzes a new PINN optimization paradigm: region optimization. Going beyond previous scatter-point optimization, we extend the optimization from selected points to their neighborhood regions. Based on this idea, RoPINN is implemented as a simple but effective training algorithm, where an effcient Monte Carlo approximation process is used along with a trust region calibration strategy to control the gradient estimation error caused by sampling, theoretically bringing a better balance of generalization and optimization. In addition to theoretical advantages, RoPINN can consistently boost the performance of various advanced PINN models without extra backpropagation or gradient calculation, demonstrating favorable efficiency, training stability and general capability.
Acknowledgments and Disclosure of Funding
This work was supported by the National Natural Science Foundation of China (U2342217 and 62022050), the BNRist Project, and the National Engineering Research Center for Big Data Software.
@@
$\begin{split}\mathbb{E}_{\mathcal{S},\mathcal{A}}\left[\mathcal{L}(\mathcal{A}(\mathcal{S}),\mathcal{S})\right]&=\mathbb{E}_{\mathcal{S},\mathcal{A}}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\mathcal{A}(\mathcal{S}),\boldsymbol{x}_{i})\right]\\ &=\mathbb{E}_{\mathcal{S},\mathcal{S}^{\prime},\mathcal{A}}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\mathcal{A}(\mathcal{S}^{(i)}),\boldsymbol{x}_{i}^{\prime})\right]\quad\quad\quad\quad\text{(Expectation on $\boldsymbol{x}_{i}^{\prime}$)}\\ &=\mathbb{E}_{\mathcal{S},\mathcal{S}^{\prime},\mathcal{A}}\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\mathcal{A}(\mathcal{S}),\boldsymbol{x}_{i}^{\prime})\right]+\delta\\ &=\mathbb{E}_{\mathcal{S},\mathcal{A}}\left[\mathcal{L}(\mathcal{A}(\mathcal{S}),\Omega)\right]+{\delta}.\end{split}$ (16)
@@
According to Lemma A.1, we attempt to bound the generalization error in expectation $\mathcal{E}_{\mathrm{gen}}$ by analyzing the error difference between two selected sample sets. Denote that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are two identical sample sets of size $|\mathcal{S}|$ except for one sample. Suppose that with the stochastic gradient method on these two sets, we can obtain two optimization trajectories $\{\theta_{t}\}_{t=1}^{T}$ and $\{\theta_{t}^{\prime}\}_{t=1}^{T}$ respectively. According to Assumption 3.2, we have the following inequality:
@@
We assume two optimization trajectories, both obtained under the same random update rule and random permutation rule. Note that at the $t$-th step, with probability $(1-\frac{1}{|\mathcal{S}|})$, the example selected by the stochastic gradient method is the same in both $\mathcal{S}$ and $\mathcal{S}^{\prime}$. As for the other $\frac{1}{|\mathcal{S}|}$ probability, we have to deal with different selected samples. Thus, according to Lemma A.2, we have:
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|\right]&=(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[\|G_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G_{\alpha_{t},\boldsymbol{x}}(\theta_{t}^{\prime})\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|G_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})\mathbb{E}[\|\theta_{t}-\theta_{t}^{\prime}\|]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+\|\alpha_{t}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x})-\alpha_{t}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}^{\prime})\|\right].\\ \end{split}$ (20)
@@
We also consider the optimization trajectory $\{\theta_{t}\}_{t=1}^{T}$ and $\{\theta_{t}^{\prime}\}_{t=1}^{T}$ from $\mathcal{S}$ and $\mathcal{S}^{\prime}$, which are identical except for one element. We assume two optimization trajectories, both obtained under the same random update rule and random permutation rule. Let $\delta_{t}=\|\theta_{t}-\theta_{t}^{\prime}\|$ and $t_{0}\in\{1,\cdots,|\mathcal{S}|\}$ be a considered iteration. Here, $t\leq|\mathcal{S}|$ because for $t>|\mathcal{S}|$, we must have $\delta_{t_{0}}\neq 0$. Then we have:
@@
Firstly, we would like to prove the convex case.